OpenStudy (kap01):

2 1/4 Divided by 3/4 simplest form

OpenStudy (anonymous):

3

OpenStudy (anonymous):

It would be- (21/4)/(3/4)= 21/3 = 7

OpenStudy (kap01):

ok thanks, but do you know how to write it out step by step?

OpenStudy (anonymous):

Is this 21/4 or 2(1/4) ???

OpenStudy (kap01):

2 (1/4)

OpenStudy (anonymous):

O.k then its going to be simplifying a mixed fraction- 2(1/4)= 9/4 then, (9/4)/ (3/4) = 9/3 = 3

OpenStudy (kap01):

thanks

OpenStudy (anonymous):

my pleasure...

OpenStudy (kap01):

what about this one? 3 (4/5) divided by 2/5

OpenStudy (anonymous):

the same procedure would do, 3(4/5) = {(3*5)+4}/5 = 19/5 Then divide the terms, as (19/5)/ (2/5) = 19/2 I hope this makes sense ...

OpenStudy (kap01):

cool thanks I think I am catching on now :)